Calculate anionic gap in following example
Na 156 mmol
K 5.6 mmol
Urea 2.9 mmol
Cl 95 mmol/L
HCO3       10 mmol
What is your diagnosis?
Answer #1

ANSWER.

* Calculation of anion gap.

        Formula = Na + K - (Cl + HCO3)

    Here

        Na = 156 mmol/l.

         K = 5.6 mmol/l.

         Cl = 95 mmol/l.

         HCO3 = 10 mmol

As per above formula

         Anion gap = 156 + 5.6 - (95+10).

                          = 161.6 - 105

                          = 56.6 mEq/L (High anion gap).

       Normal range of anion gap = 12 - 16 mEq/L (When potassium included).

* Diagnosis : High anion gap metabolic acidosis.

                        (Presence of increased amount of acid in patients blood).